A proton travels in the accelerator at Fermilab at 99.999954% of the speed of light. An anti-proton travels in the opposite direction at the same speed. What is the relative velocity of the two particles?
Answer #1

let

u = 99.999954 % of C

= 0.99999954*c

v = 99.999954 % of C

= 0.99999954*c

relative velocity of the two particles = ( u + v)/(1 + u*v/c^2)

= (0.99999954 + 0.99999954)/(1 + 0.99999954*c*0.99999954*c/c^2)

= 0.9999999999999*c <<<<<----Answer

= 0.9999999999999*3*10^8 m/s

= 2.99999999.99997002 m/s